# Learning Exactly Linearizable Deep Dynamics Models

### v1

@@ 1 Introduction
In recent years, there has been a growing interest in using machine learning (particularly deep learning) for modeling dynamical systems (Kocijan et al., 2004; Hedjar, 2013; Lenz et al., 2015; Moriyasu et al., 2019). Unlike the traditional modeling approach, machine learning can be used to create highly accurate models more easily without the need for detailed domain knowledge. However, control design using such models can become challenging because of the strong nonlinearity of the machine-learning models. Model predictive control (MPC) is a typical control method that can handle nonlinearity, wherein the model is used to optimize the predicted future behavior of the system; however, the optimal control problem solved each time is often nonconvex (Ławryńczuk, 2008; Nghiem, 2019; Gros, 2019) when the model is highly nonlinear, thereby making it difficult to ensure the uniqueness and optimality of the numerically obtained solution and the continuity of the control law (Moriyasu et al., 2022). Various nonlinear control theories besides MPC cannot be easily applied to general machine learning models because the model structures to which the theories can be applied are often limited to specific systems, such as input-affine systems. Machine learning models with special structures for controlling design methods have been proposed to overcome the abovementioned problem. For example, a dynamical model structure called the input convex recurrent neural network (ICRNN) (Chen et al., 2019) was proposed to guarantee the convexity of the optimal control problem in economic MPC, i.e., MPC that directly aims to minimize or maximize the outputs or states. This is an extension of the idea of an input convex neural network (ICNN) (Amos et al., 2017), which is a deep neural network that can guarantee the convexity of the input–output relationships, to dynamical systems. Further, a structured Hammerstein–Wiener (S–HW) model (Moriyasu et al., 2022) can ensure the convexity of the usual MPC problems for reference tracking or regulation for which convexity cannot be guaranteed with ICRNN. This is an application of the ICNN and bijective neural network (BNN) (Baird et al., 2005) to the Hammerstein–Wiener (HW) model (Hammerstein, 1930; Wiener, 1958), which has long been known in the field of system identification. The HW model has a linear dynamical system sandwiched between static nonlinear bijective mappings. Static nonlinearity can be canceled out by utilizing bijectivity, and the control problems can be reduced to linear control problems (Fruzzetti et al., 1997; Cervantes et al., 2003). In the S–HW model, BNNs are used to express bijective mappings and an ICNN is attached to learn the additional outputs to be constrained. Therefore, the constrained nonlinear tracking control problems can be reduced to convex problems. However, such a specific structured model can lose its expressive ability when the above properties are ensured by limiting the range of weights, choice of activation function, and connecting topology of the network. Therefore, it is necessary to propose a model structure that can learn a wider range of objects with high accuracy and is easy to use in control design. Given this context, we focused on a class of nonlinear systems called exactly linearizable (EL) systems and proposed a learning model that is guaranteed to have this property. Although the HW model represents the input/output of a nonlinear system by transforming the input/output of a linear dynamical system using static nonlinear functions, EL systems can be attributed to linear dynamics using nonlinear dynamic feedback (Khalil, 2002). This feedback is an extension of static mapping, and therefore, an EL system can be considered an extension of the HW system. The EL models lose the ability to make the constrained MPC convex because of the nonlinear feedback structure. However, because they are linearizable, linear control methods, such as linear quadratic regulators (LQR), and various nonlinear control methods can be easily applied with control barrier functions (CBF), which enables the constraint-aware control design. The contribution of this study is that it proposes an approach to learn EL models, which is a class of systems that have a higher degree of freedom of expression than conventional methods, while facilitating control design. Verifying that a dynamical system is EL and constructing a dynamic feedback law that achieves linearization is usually based on heuristics and is therefore impractical. Our idea of introducing output feedback in a static nonlinear mapping of the S–HW model structure can guarantee that the learning model has EL properties and can explicitly obtain a dynamic feedback law. The remainder of this paper is organized as follows: Section 2 shows the structure of the proposed model in comparison with the S–HW model. Section 3 presents a control design method using the proposed model. Further, we present the simplest method based on LQR. Further, this section introduces integral CBFs (Ames et al., 2021) which enable the model to be more flexible in terms of its degrees of freedom. In Section 4, the model is applied to the air path and combustion process of an engine, which is a complex nonlinear system, to verify its advantages in terms of accuracy and show that control objectives can be successfully achieved via a control simulation.
@@ 2.1 Problem Statement
This study addresses the modeling of nonlinear MIMO systems and the design of control based on the obtained models considering the upper and lower input constraints and the upper output constraint. Let $v\in\mathbb{R}^{m}$, $d\in\mathbb{R}^{l}$, $y\in\mathbb{R}^{n}$, and $z\in\mathbb{R}^{p}$ represent the input, disturbance, output for tracking control, and constrained output, respectively. We consider building a model that can predict the response $y,z$ to $v,d$ by deep learning. In control design, the controller is designed to maintain $y$ close to the target $y_{\mathrm{d}}\in\mathbb{R}^{n}$ considering the upper and lower constraints for each element of the input $v=[v_{1},\ldots,v_{m}]^{\mathrm{T}}$
$\underline{v}_{i}\leq v_{i}\leq\overline{v}_{i}\ (i=1,\ldots,m)$ (1)
and the upper constraint for each element on the constraint output $z=[z_{1},\ldots,z_{p}]^{\mathrm{T}}$
$z_{j}\leq\overline{z}_{j}\ (j=1,\ldots,p).$ (2)
The disturbance $d$ cannot be adjusted; however, it can be observed online.
@@ 2.2 Exact Linearization
$\dot{y}=\overline{f}(y)+\overline{g}(y)v$ (3)
with $y\in\mathbb{R}^{n},v\in\mathbb{R}^{m}$. In the exact linearization method, system (3) is linearized using nonlinear feedback and coordinate transformation, which are denoted by
$x$ $=\Phi(y)$ (4)
$v$ $=\Psi_{1}(y)+\Psi_{2}(y)u$ (5)
using a bijective map $\Phi$. This yields
$\dot{x}$ $=\frac{\partial\Phi}{\partial y}\{\overline{f}(\Phi^{-1}(x))+\overline{g}(\Phi^{-1}(x))\Psi_{1}(\Phi^{-1}(x))\}+\frac{\partial\Phi}{\partial y}\overline{g}(\Phi^{-1}(x))\Psi_{2}(\Phi^{-1}(x))u$
$=:f(x)+g(x)u.$ (6)
@@
$f(x)=Ax,\ g(x)=B$ (7)
and $(A,B)$ are controllable, the system is considered exactly linearizable. Assuming that the plant and feedback are input-affine, as in (3) and (5), such a linearizing transformation and feedback exist under some theoretically sufficient conditions; see Appendix 1. However, this input-affine structure is excessively restrictive for the dynamics of interest, and therefore, we assumed the following target system, coordinate transformation, and feedback:
$\dot{y}=\overline{f}(v,y),\ x=\Phi(y),\ v=\Psi(u,y)$ (8)
for which the dynamics of $x$ are affine such that
$\dot{x}=$ $Ax+Bu+c.$ (9)
@@
Suppose $\Psi:\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{l}\rightarrow\mathbb{R}^{m}$, $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{l}\rightarrow\mathbb{R}^{n}$ and $\Xi:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{l}\rightarrow\mathbb{R}^{p}$ satisfy the following conditions: $\Psi(\cdot,y,d),\ \Phi(\cdot,d)$ are bijective. $\Xi(\cdot,\cdot,d)$ is convex for any $y,d$, then, the system
$u(t)$ $=\Psi^{-1}(v(t),y(t),d(t)),$ (10)
$\dot{x}(t)$ $=A(d(t))x(t)+B(d(t))u(t)+c(d(t)),$ (11)
$y(t)$ $=\Phi^{-1}(x(t),d(t)),$ (12)
$z(t)$ $=\Xi(x(t),u(t),d(t))$ (13)
is denoted as an exactly linearizable model, where $t\in\mathbb{R}$ and $u\in\mathbb{R}^{m}$ and $x\in\mathbb{R}^{n}$ represent the continuous time and the internal input and state of the affine dynamics, respectively.
@@ 2.3 Exactly Linearizable Model
In this section, we propose a model structure that corresponds to the Definition 1, which we call the EL model. A block diagram of the EL model is shown in Fig. 1. The red line in Fig. 1 shows its difference from the HW model (Moriyasu et al., 2022).
@@
The HW model refers to the model in which (10) is restricted to the output-independent $u(t)=\Psi^{-1}(v(t),d(t))$. This simple extension allows the model to learn a wider range of dynamical systems than the S–HW model. Further, the S–HW model always has a unique equilibrium corresponding to each input, whereas the EL model can represent systems with multiple equilibria. This difference greatly improves the accuracy of the model but requires more careful treatment during control design. We attempt to construct EL models introduced in Definition 1 from an input–output data sequence. Recall that functions $\Phi(\cdot,d)$ and $\Psi(\cdot,y,d)$ must be bijective mappings. We parameterize each mapping using a bijective neural network (BNN) (Baird et al., 2005) and a diagonal-BNN, respectively. A BNN represents bijective mapping as the $N$-layered neural network $\Phi(y,d)=\varphi^{(N)}\circ\varphi^{(N-1)}\circ\cdots\circ\varphi^{(1)}(y,d)$, where $\varphi^{(i)}(\cdot,d)(i=1,\ldots,N)$ are parametrized bijective mappings. For example, $\varphi^{(i)}(y,d)=\sinh^{-1}(c^{(i)}(d)+\sinh(W^{(i)}(d)y+b^{i}(d)))$, where $W^{(i)}(d)$ and $b^{(i)}(d),c^{(i)}(d)$ represents matrix-valued and vector-valued functions to be learned. If $W^{(i)}(d)$ is diagonal for all $d\in\mathbb{R}^{l}$ and $i=1,\cdots,N$, the input–output relationship of the BNN becomes element-wise, which we call the diagonal-BNN. Moreover, the function $\Xi(\cdot,\cdot,d)$ must be convex, and we adopt a partially input convex neural network (PICNN) (Amos et al., 2017) for parametrization. The structure of the PICNN is rather complex, and therefore, it is omitted in this paper for simplicity. This model includes many disturbance-dependent parameters such as $A,B,c,W^{(i)},b^{(i)},c^{(i)}(i=1,\cdots,N)$. All these can be arbitrary functions; however, in this study, a standard three-layer fully connected neural network is used to represent each element. A summary of the superior properties of the EL model is presented below. First, the dynamical system can be considered linear because of the dynamic feedback. Second, the reference for $y$ can be replaced with that for $x$ because of the bijectivity of $\Phi$. Third, the admissible set of $u$ corresponding to the input constraint (1) is convex for any $x$ owing to the element-wise bijectivity of $\Psi$. Finally, the admissible set of $x,u$ corresponding to the output constraint (2) is convex because of the convexity of $\Xi$. The learning process of the EL model is as follows: The EL model structure is similar to that of the S–HW model (Moriyasu et al., 2022), except that output $y$ is fed back into (10), to use the same procedure for learning the model. The outputs predicted by the model are expressed with a hat to distinguish it from real data. From (10)–(13), we obtain
$\dot{\hat{y}}$ $={\left(\frac{\partial\Phi}{\partial y}\right)}^{-1}{\left(A(d)\Phi(y,d)+B(d)\Psi^{-1}(v,y,d)+c(d)-\frac{\partial\Phi}{\partial d}\dot{d}\right)},$ (14)
$\hat{z}$ $=\Xi(\Phi(y,d),\Psi^{-1}(v,y,d),d).$ (15)
@@
A key point is that the bijectivity of the BNN enables these evaluations without using internal signals $x,u$. The data required to evaluate these are those of $v,y,d,\dot{d}$. The prediction error can be evaluated as $e:=[(\dot{\hat{y}}-\dot{y})^{\mathrm{T}},(\hat{z}-z)^{\mathrm{T}}]^{\mathrm{T}}$ using the answer data of $\dot{y},z$. Thus, model learning is reduced to an error minimization problem
@@
where $N_{s}$, $e_{i}$, $Q_{e}$, and $\theta$ represent the number of sampled data points, prediction error for the $i$ th data point, positive definite matrix, and model parameter vector, respectively.

### v2

@@ Abstract
Keywords: model predictive control, machine learning, control application, nonlinear control, Hammerstein-Wiener model
@@ 1 Introduction
In recent years, there has been a growing interest in using machine learning (particularly deep learning) for modeling dynamical systems (Kocijan et al., 2004; Hedjar, 2013; Lenz et al., 2015; Moriyasu et al., 2019). Unlike the traditional modeling approach, machine learning can be used to create highly accurate models more easily without the need for detailed domain knowledge. However, control design using such models can become challenging because of the strong nonlinearity of the machine-learning models. Model predictive control (MPC) is a typical control method that can handle nonlinearity, wherein the model is used to optimize the predicted future behavior of the system; however, the optimal control problem solved each time is often nonconvex (Ławryńczuk, 2008; Nghiem, 2019; Gros, 2019) when the model is highly nonlinear, thereby making it difficult to ensure the uniqueness and optimality of the numerically obtained solution and the continuity of the control law (Moriyasu et al., 2022). Various nonlinear control theories besides MPC cannot be easily applied to general machine learning models because the model structures to which the theories can be applied are often limited to specific systems, such as input-affine systems.
Machine learning models with special structures for controlling design methods have been proposed to overcome the abovementioned problem. For example, a dynamical model structure called the input convex recurrent neural network (ICRNN) (Chen et al., 2019) was proposed to guarantee the convexity of the optimal control problem in economic MPC, i.e., MPC that directly aims to minimize or maximize the outputs or states. This is an extension of the idea of an input convex neural network (ICNN) (Amos et al., 2017), which is a deep neural network that can guarantee the convexity of the input–output relationships, to dynamical systems. Further, a structured Hammerstein–Wiener (S–HW) model (Moriyasu et al., 2022) can ensure the convexity of the usual MPC problems for reference tracking or regulation for which convexity cannot be guaranteed with ICRNN. This is an application of the ICNN and bijective neural network (BNN) (Baird et al., 2005) to the Hammerstein–Wiener (HW) model (Hammerstein, 1930; Wiener, 1958), which has long been known in the field of system identification. The HW model has a linear dynamical system sandwiched between static nonlinear bijective mappings. Static nonlinearity can be canceled out by utilizing bijectivity, and the control problems can be reduced to linear control problems (Fruzzetti et al., 1997; Cervantes et al., 2003). In the S–HW model, BNNs are used to express bijective mappings and an ICNN is attached to learn the additional outputs to be constrained. Therefore, the constrained nonlinear tracking control problems can be reduced to convex problems.
However, such a specific structured model can lose its expressive ability when the above properties are ensured by limiting the range of weights, choice of activation function, and connecting topology of the network. Therefore, it is necessary to propose a model structure that can learn a wider range of objects with high accuracy and is easy to use in control design.
Given this context, we focused on a class of nonlinear systems called exactly linearizable (EL) systems and proposed a learning model that is guaranteed to have this property. Although the HW model represents the input/output of a nonlinear system by transforming the input/output of a linear dynamical system using static nonlinear functions, EL systems can be attributed to linear dynamics using nonlinear dynamic feedback (Khalil, 2002). This feedback is an extension of static mapping, and therefore, an EL system can be considered an extension of the HW system. The EL models lose the ability to make the constrained MPC convex because of the nonlinear feedback structure. However, because they are linearizable, linear control methods, such as linear quadratic regulators (LQR), and various nonlinear control methods can be easily applied with control barrier functions (CBF), which enables the constraint-aware control design.
The contribution of this study is that it proposes an approach to learn EL models, which is a class of systems that have a higher degree of freedom of expression than conventional methods, while facilitating control design. Verifying that a dynamical system is EL and constructing a dynamic feedback law that achieves linearization is usually based on heuristics and is therefore impractical. Our idea of introducing output feedback in a static nonlinear mapping of the S–HW model structure can guarantee that the learning model has EL properties and can explicitly obtain a dynamic feedback law.
The remainder of this paper is organized as follows: Section 2 shows the structure of the proposed model in comparison with the S–HW model. Section 3 presents a control design method using the proposed model. Further, we present the simplest method based on LQR. Further, this section introduces integral CBFs (Ames et al., 2021) which enable the model to be more flexible in terms of its degrees of freedom. In Section 4, the model is applied to the air path and combustion process of an engine, which is a complex nonlinear system, to verify its advantages in terms of accuracy and show that control objectives can be successfully achieved via a control simulation.
@@ 2.1 Problem Statement
This study addresses the modeling of nonlinear MIMO systems and the design of control based on the obtained models considering the upper and lower input constraints and the upper output constraint. Let $u\in\mathbb{R}^{m}$, $d\in\mathbb{R}^{l}$, $x\in\mathbb{R}^{n}$, and $y\in\mathbb{R}^{p}$ represent the control input, exogenous input (which can include measurable disturbance), state, and output, respectively. We assume that the state $x$, exogenous input $d$, and their time-derivatives $\dot{x},\dot{d}$ (numerical differentiation is acceptable) can be observed online and $d$ cannot be adjusted. Observation noise is not considered explicitly in this paper.
We consider building a model that can predict the response $x,y$ to $u,d$ by deep learning. In control design, the controller is designed to maintain $x$ close to the target $x_{\mathrm{d}}\in\mathbb{R}^{n}$ considering the upper and lower constraints for each element of the control input $u=[u_{1},\ldots,u_{m}]^{\mathrm{T}}$
$\underline{u}_{i}\leq u_{i}\leq\overline{u}_{i}\ (i=1,\ldots,m)$ (1)
and the upper constraint for each element on the output $y=[y_{1},\ldots,y_{p}]^{\mathrm{T}}$
$y_{j}\leq\overline{y}_{j}\ (j=1,\ldots,p).$ (2)
@@ 2.2 Exact Linearization
$\dot{x}=\overline{f}(x)+\overline{g}(x)u$ (3)
with $x\in\mathbb{R}^{n},u\in\mathbb{R}^{m}$. In the exact linearization method, system (3) is transformed by using coordinate transformation and nonlinear feedback, which are denoted by
$\xi$ $=\Phi(x)$ (4)
$u$ $=\Psi_{1}(x)+\Psi_{2}(x)v$ (5)
with a bijective map $\Phi$, nonlinear functions $\Psi_{1},\Psi_{2}$, and newly introduced input $v\in\mathbb{R}^{m}$. This yields
$\dot{\xi}$ $=\frac{\partial\Phi}{\partial x}\{\overline{f}(\Phi^{-1}(\xi))+\overline{g}(\Phi^{-1}(\xi))\Psi_{1}(\Phi^{-1}(\xi))\}+\frac{\partial\Phi}{\partial x}\overline{g}(\Phi^{-1}(\xi))\Psi_{2}(\Phi^{-1}(\xi))v$
$=:f(\xi)+g(\xi)v.$ (6)
@@
$f(\xi)=A\xi,\ g(\xi)=B$ (7)
and $(A,B)$ are controllable, the system is considered exactly linearizable.
Assuming that the plant and feedback are input-affine, as in (3) and (5), such a linearizing transformation and feedback exist under some theoretically sufficient conditions; see Appendix 1. However, this input-affine structure is sometimes restrictive for the dynamics of complex real systems, and therefore, we assumed the following target system, coordinate transformation, and feedback:
$\dot{x}=\overline{f}(x,u),\ \xi=\Phi(x),\ u=\Psi^{-1}(v,x)$ (8)
for which the dynamics of $\xi$ are affine such that
$\dot{\xi}=$ $A\xi+Bv+c.$ (9)
@@
Suppose $\Psi:\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{l}\rightarrow\mathbb{R}^{m}$, $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{l}\rightarrow\mathbb{R}^{n}$ and $\Xi:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{l}\rightarrow\mathbb{R}^{p}$ satisfy the following conditions: $\Psi(\cdot,x,d),\ \Phi(\cdot,d)$ are bijective. $\Xi(\cdot,\cdot,d)$ is convex for any $x,d$, then, the system
$v(t)$ $=\Psi(u(t),x(t),d(t)),$ (10)
$\dot{\xi}(t)$ $=A(d(t))\xi(t)+B(d(t))v(t)+c(d(t)),$ (11)
$x(t)$ $=\Phi^{-1}(\xi(t),d(t)),$ (12)
$y(t)$ $=\Xi(\xi(t),v(t),d(t))$ (13)
is denoted as an exactly linearizable model, where $t\in\mathbb{R}$ and $v\in\mathbb{R}^{m}$ and $\xi\in\mathbb{R}^{n}$ represent the continuous time and the internal input and state of the affine dynamics, respectively.
Note that the drift term $c$ in (9) and (11) is introduced to extend the degree of freedom of the model and is removable if it is unnecessary.
@@ 2.3 Exactly Linearizable Model
In this section, we propose a neural network based model structure that corresponds to the Definition 1, which we call the EL model. A block diagram of the EL model is shown in Fig. 1. The red line in Fig. 1 shows its difference from the S–HW model (Moriyasu et al., 2022). The S–HW model refers to the model in which (10) is restricted to the output-independent $v(t)=\Psi(u(t),d(t))$. This simple extension in the EL model allows the model to learn a wider range of dynamical systems than the S–HW model. For example, the S–HW model always has a unique equilibrium corresponding to each input, whereas the EL model can represent systems with multiple equilibria. This difference greatly improves the accuracy of the model but requires more careful treatment during control design.
@@
Figure 2: Network structure of exactly linearizable model
The whole network structure of EL models is illustrated in Fig. 2. Recall that functions $\Phi(\cdot,d)$ and $\Psi(\cdot,x,d)$ must be bijective mappings. We parameterize each mapping using a bijective neural network (BNN) (Baird et al., 2005) and a diagonal-BNN, respectively. $\Phi$ is represented as the $N_{\Phi}$-layered BNN
$\Phi(x,d)=\varphi^{(N_{\Phi})}_{\Phi}\circ\varphi^{(N_{\Phi}-1)}_{\Phi}\circ\cdots\circ\varphi^{(1)}_{\Phi}(x,d),$ (14)
where $\varphi^{(i)}_{\Phi}(\cdot,d)(i=1,\ldots,N_{\Phi})$ are parametrized bijective mappings. In this study, we employ
$\varphi^{(i)}_{\Phi}(x,d)=\sinh^{-1}(a^{(i)}_{\Phi}(d)+\sinh(W^{(i)}_{\Phi}(d)x+b^{(i)}_{\Phi}(d)))\ (i=1,\ldots,N_{\Phi}),$ (15)
where $W^{(i)}_{\Phi}(d)$ and $a^{(i)}_{\Phi}(d),b^{(i)}_{\Phi}(d)$ represents matrix-valued and vector-valued functions to be learned, described later. Similarly, $\Psi$ is represented as the $N_{\Psi}$-layered diagonal BNN
$\Psi(u,x,d)$ $=\varphi^{(N_{\Psi})}_{\Psi}\circ\varphi^{(N_{\Psi}-1)}_{\Psi}\circ\cdots\circ\varphi^{(1)}_{\Psi}(u,x,d),$ (16)
$\varphi^{(i)}_{\Psi}(u,x,d)$ $=\sinh^{-1}(a^{(i)}_{\Psi}(x,d)+\sinh(W^{(i)}_{\Psi}(x,d)u+b^{(i)}_{\Psi}(x,d)))\ (i=1,\ldots,N_{\Psi}).$ (17)
Unlike $\Phi$, functions $W^{(i)}_{\Psi}\ (n=1,\cdots,N_{\Psi})$ have to be diagonal so that the input–output relationship of the BNN becomes element-wise. Moreover, the function $\Xi(\cdot,\cdot,d)$ must be convex, and we adopt a partially input convex neural network (PICNN) (Amos et al., 2017) for parametrization. PICNN is represented as the $N_{\Xi}$-layered neural network:
$\Xi(\xi,v,d)$ $=\zeta^{(N_{\Xi})},\zeta^{(0)}=[\xi^{\mathrm{T}},v^{\mathrm{T}}]^{\mathrm{T}},\eta^{(0)}=d,$ (18)
$\zeta^{(i)}$ $=\varphi_{\Xi}^{(i)}\left(W_{\zeta}^{(i)}{\left(\zeta^{(i-1)}\odot{\rm softplus}(W_{\zeta\eta}^{(i)}\eta^{(i-1)}+b_{\zeta\eta}^{(i)})\right)}\right.$
$\ \ \ \ \ \ \ \ \ \left.+\ W_{0}^{(i)}{\left(\zeta^{(0)}\odot(W_{0\eta}^{(i)}\eta^{(i-1)}+b_{0\eta}^{(i)})\right)}+W_{b\eta}^{(i)}\eta^{(i-1)}+b_{b\eta}^{(i)}\right),$ (19)
$\eta^{(i)}$ $=\varphi_{\Xi d}^{(i)}{\left(W_{\eta}^{(i)}\eta^{(i-1)}+b_{\eta}^{(i)}\right)},(i=1,\ldots,N_{\Xi}),$ (20)
where $W_{\zeta}^{(i)},W_{0}^{(i)},W_{\eta}^{(i)},W_{\zeta\eta}^{(i)},W_{0\eta}^{(i)},W_{b\eta}^{(i)}\ (i=1,\ldots,N_{\Xi})$ and $b_{\eta}^{(i)},b_{\zeta\eta}^{(i)},b_{0\eta}^{(i)},b_{b\eta}^{(i)}\ (i=1,\ldots,N_{\Xi})$ are weights and biases to be learned, $\varphi_{\Xi}^{(i)},\varphi_{\Xi d}^{(i)}\ (i=1,\ldots,N_{\Xi})$ are activation functions, and ${\rm softplus}(x):=\log(1+e^{x})$. $\varphi_{\Xi}^{(i)},W_{\zeta}^{(i)}\ (i=1,\ldots,N_{\Xi})$ have to be monotonically non-decreasing convex functions and nonnegative matrices, respectively, to ensure convexity of $\Xi$ with respect to $\xi,v$.
In this model, $d$-dependent functions $A,B,c,W^{(i)}_{\Phi},a^{(i)}_{\Phi},b^{(i)}_{\Phi},W^{(j)}_{\Psi},a^{(j)}_{\Psi},b^{(j)}_{\Psi}\ (i=1,\ldots,N_{\Phi},\ j=1,\ldots,N_{\Psi})$ can be arbitrary functions; however, in this study, a standard fully connected neural network (FNN) is used to represent each element. Precisely, $W^{(j)}_{\Psi},a^{(j)}_{\Psi},b^{(j)}_{\Psi}\ (j=1,\ldots,N_{\Psi})$ also have $x$-dependency in EL model. We deal with this $(x,d)$-dependency by employing $x$-dependent FNN and add its output with that of $d$-dependent FNN to obtain the above functions (shown as red in Fig. 2). The example of detailed settings is show in Section 4.
A summary of the superior properties of the EL model is presented below. First, the dynamical system can be considered linear because of the dynamic feedback. Second, the reference for $x$ can be replaced with that for $\xi$ because of the bijectivity of $\Phi$. Third, the admissible set of $v$ corresponding to the input constraint (1) is convex for any $\xi$ owing to the element-wise bijectivity of $\Psi$. Finally, the admissible set of $\xi,v$ corresponding to the output constraint (2) is convex because of the convexity of $\Xi$.
The learning process of the EL model is as follows: The EL model structure is similar to that of the S–HW model (Moriyasu et al., 2022), except that output $x$ is fed back into (10), to use the same procedure for learning the model. The outputs predicted by the model are expressed with a hat to distinguish it from real data. From (10)–(13), we obtain
$\dot{\hat{x}}$ $={\left(\frac{\partial\Phi}{\partial x}\right)}^{-1}{\left(A(d)\Phi(x,d)+B(d)\Psi(u,x,d)+c(d)-\frac{\partial\Phi}{\partial d}\dot{d}\right)},$ (21)
$\hat{y}$ $=\Xi(\Phi(x,d),\Psi(u,x,d),d).$ (22)
@@
A key point is that the bijectivity of the BNN enables these evaluations without using internal signals $\xi,v$. The data required to evaluate these are those of $u,x,d,\dot{d}$. The prediction error can be evaluated as $e:=[(\dot{\hat{x}}-\dot{x})^{\mathrm{T}},(\hat{y}-y)^{\mathrm{T}}]^{\mathrm{T}}$ using the answer data of $\dot{x},y$. Thus, model learning is reduced to an error minimization problem
@@
where $N_{s}$, $e_{i}$, $Q_{e}$, and $\theta$ represent the number of sampled data points, prediction error for the $i$ th data point, positive definite matrix, and model parameter vector, respectively. $\theta$ includes all free parameters, i.e., weight matrices ans bias vectors, of the network (Fig. 2) and is not listed here due to the sheer number of elements. The above problem can be solved typically by stochastic gradient descent (SGD) with a machine learning framework such as Tensorflow. Using such tools, Jacobians $\partial\Phi/\partial x,\partial\Phi/\partial d$ and also their derivatives, which are required in gradient based learning algorithms, can be caluculated automatically.
Note that, (22) is used only in learning phase. In control computation, we can use (11) and (12) to evaluate output $y$. That allows us to avoid heavy computation for evaluating Jacobian and achieve fast computation of model-based controller. In addition, data, especially of time derivatives ($\dot{x},\dot{d}$), are often noisy in practice, but in deep learning, a certain amount of noise may be rather beneficial to improve generalization performance (Liu et al., 2020). If simulators or very accurate experimental instruments are used to generate data, we recommend adding some noise to avoid overfitting.